OpenStudy (anonymous):

I have the linear equations: -4x-7y=5 and -3x-2y=11. I need to solve for x and y.

OpenStudy (texaschic101):

substitution or elimination ?

OpenStudy (mrnood):

multiply the first equation by 3 multiply the second equation by 4 subtract the 2 new equations This will give you an equqtion in y only - solve for y puyt this VALUE of y into one of the original equations - this will give an equation in x only - solve for x...

OpenStudy (texaschic101):

-4x - 7y = 5 ---->(2) -3x - 2y = 11 --->(-7) -------------------- -8x - 14y = 10 (result of multiplying by 2) 21x + 14y = -77 (result of multiplying by -7) --------------------add 13x = - 67 x = -67/13 now sub -67/13 in for x in either equation to find y -3x - 2y = 11 -3(-67/13) - 2y = 11 201/13 - 2y = 11 -2y = 11 - 201/13 -2y = 143/13 - 201/13 -2y = - 58/13 y = (-58/13) / (-2) y = -58/13 * - 1/2 y = 58/26 which reduces to 29/13 check... -3x - 2y = 11 -3(-67/13) - 2(29/13) = 11 201/13 - 58/13 = 11 143/13 = 11 11 = 11 (correct) so, x = -67/13 and y = 29/13.....or (-67/13, 29/13)